OpenStudy (anonymous):

*Algebraic fraction* I dont need the answer just a walk though! so confused :( I am part way though 1/x-2 + x-3 / (x+1)(x-2)

OpenStudy (anonymous):

ok try to take the common denominator (x+1)(x-2) and do the other steps

OpenStudy (anonymous):

[(x+1) +(x^2-5x+6)]/ (x+1)(x-2)

OpenStudy (anonymous):

= (x^2-4x+7)/(x^2-x-2)

OpenStudy (anonymous):

Sorry I really dont know how you got there?! I have got 1(x-1) + (x-3) / (x+1) (x-2)

OpenStudy (anonymous):

I am trying to simplfy! And I have got an asnwer! I just cant get to it. 2(x-1) / (x-2)(x-1)

OpenStudy (anonymous):

well. let's check again

OpenStudy (anonymous):



OpenStudy (anonymous):

Multiply the first fraction by x+1 over itself to get an lcd, then add the fractions together\[\frac{1}{x-2}+\frac{x-3}{(x+1)(x-2)}=\frac{x+1+x-3}{(x+1)(x-2)}=\frac{2x-2}{(x+1)(x-2)}\]you could multiply the denominator, which is considered more simplified\[\frac{2x-2}{x^2-x-2}\]

OpenStudy (anonymous):

then expand the numerator & denominator

OpenStudy (anonymous):

@jpk, I assumed that both binomials in the second fraction are in the denominator. is that correct?

OpenStudy (anonymous):

@man they were already like that so I assume soo. Ah! I see I got my signs mixed up! :( Thanks guys! It helpd!